# Vendor-Aware Industrial Agents: RAG-Enhanced LLMs for Secure On-Premise PLC Code Generation

### v1

@@ I Introduction
Every day, millions of Programmable Logic Controllers (PLCs) orchestrate production lines, manage power grids, and safeguard critical infrastructure—from automotive assembly plants to water treatment facilities [John2010]. Behind each automated process lies control logic programmed in specialized languages, most prominently Structured Text (ST), a high-level language standardized in IEC 61131-3 that combines the expressiveness of modern programming with the real-time determinism industrial systems demand [John2010].
@@
Despite ST’s syntactic similarity to languages like Python or Pascal, developing PLC control logic remains fundamentally different. Where Python developers access thousands of open-source libraries and Stack Overflow answers, ST engineers face sparse code repositories and proprietary vendor documentation. Moreover, each manufacturer – Siemens, CODESYS, Mitsubishi Electric, Rockwell Automation – implements distinct dialects with incompatible function blocks, naming conventions, and compiler behaviors [Yang2024, Haag2024]. A timer implementation for Siemens TIA Portal will not compile on a Mitsubishi Electric iQ-R controller; even seemingly portable constructs like counter instructions exhibit vendor-specific reset semantics and enable/disable conventions.
@@
This fragmentation collides with industrial realities: manufacturing firms guard control logic as trade secrets, regulatory frameworks restrict proprietary code sharing, and security-critical infrastructure operators prohibit cloud-dependent AI tools that could expose operational intelligence [Brehme2025]. The result is a low-data domain where traditional large language model training approaches—reliant on vast public code corpora—struggle to generate vendor-compliant, compilable control logic [Haag2024]. Existing coding assistants trained on GitHub’s predominantly web and enterprise software fail to recognize that RTRIG_P triggers on rising edges in Mitsubishi Electric ST, that variable declarations must occur in external label editors rather than inline VAR blocks, or that certain ladder logic instructions are forbidden in ST contexts.
@@
Nonetheless, recent research has demonstrated significant advances in automated ST code generation using Large Language Models (LLMs) [Xia2025]. Haag et al. [Haag2024] pioneered the integration of compiler feedback with LLM training, achieving 70% compilation success rates through iterative Direct Preference Optimization (DPO). Liu et al. [Liu2024] developed Agents4PLC, achieving up to 68.8% formal verification success across 23 programming tasks. Yang et al. [Yang2024] advanced the field with AutoPLC, a framework for vendor-aware ST code generation achieving 60% compilation success across Siemens TIA Portal and CODESYS platforms. Meanwhile, Ren et al. [Ren2025] introduced MetaIndux-PLC, employing control logic-guided iterative fine-tuning.
@@
However, despite these promising academic results, significant gaps remain between research achievements and industrial deployment requirements [Brehme2025]. Brehme et al.’s [Brehme2025] comprehensive industry study revealed that most RAG applications in industrial settings remain in prototype stages, with data protection, security, and quality identified as paramount concerns (8.9, 8.5, and 8.7 of 10 points).
@@
Existing research has primarily focused on either general IEC 61131-3 compatibility or multiple vendor platforms simultaneously, but has not adequately addressed the challenge of creating practical coding assistants for specific proprietary implementations [Brehme2025]. While Haag et al. [Haag2024] demonstrated compiler feedback integration, their approach requires extensive fine-tuning and relies on cloud-based services. Yang et al. [Yang2024] addressed multi-vendor variations but depends on cloud-based processing incompatible with industrial security requirements [Brehme2025]. Liu et al. [Liu2024] operates as a single-model solution lacking comprehensive retrieval mechanisms for proprietary function blocks. Ren et al. [Ren2025] requires substantial fine-tuning datasets typically unavailable in proprietary environments.
@@
Recent work has extended LLM-based control logic generation beyond textual notations. Koziolek et al. [Koziolek2025] demonstrated Spec2Control, a workflow generating graphical DCS control logic for ABB systems with a high connection accuracy across 65 test cases, achieving 94-96% labor savings. While Spec2Control targets multi-plant DCS graphical programming, our work focuses on vendor-specific textual ST generation.
@@
The specialized nature of industrial automation creates a low-data domain scenario where comprehensive training datasets for specific vendor implementations are scarce or confidential [Haag2024], making traditional fine-tuning approaches impractical [Brehme2025]. Notably, no existing work has specifically addressed Mitsubishi Electric’s ST implementation and MELSOFT GX Works3, the engineering environment for Mitsubishi Electric’s PLC series, which presents challenges due to its proprietary function blocks, specific syntax variations, and integration requirements. These are, however, compliant with the IEC 61131-3 standard, since it leaves freedoms to vendors. These freedoms and said function blocks, are not known to the wider public and hence not known to LLMs.
@@
This work addresses this critical gap through a novel RAG-enhanced coding assistant specifically designed for Mitsubishi Electric’s ST implementation with the development environment and compiler MELSOFT GX Works3. Unlike previous approaches targeting multiple vendors [Yang2024] or require extensive fine-tuning [Haag2024, Ren2025], our solution focuses on one ecosystem.
@@
The key contributions include: (1) the first RAG-based coding assistant specifically designed for Mitsubishi Electric’s ST implementation, addressing a significant gap in vendor-specific industrial AI tools; (2) novel competitive multi-model approach with MELSOFT GX Works3 integration enabling real-time compilation feedback and error correction; (3) comprehensive evaluation using Mitsubishi Electric’s development environment with both compilation statistics and expert assessments from engineers familiar with the platform; and (4) practical demonstration that RAG-supported coding assistants can achieve high-quality code generation for specific proprietary platforms through targeted knowledge base construction and vendor-specific prompt engineering. We also enhance earlier work [Reinhardt2025] that used translation transformers due to privacy constraints and plan to further support local Small and Medium Sized companies (SMEs) in the future with our results. Fig. 1 illustrates an ST program generated by our approach.
@@
Figure 1: Example of generated ST code using our GeneriST web app (built with Streamlit [streamlit2025]).
@@ II Industrial Project Background and System Architecture
GX Works3 emphasizes a label- and device-based model in which variables (labels) are defined and managed in the engineering tool and then referenced by ST programs; the operating manual details dedicated editors and the standard workflow for creating ST programs, running syntax checks, and converting projects [MEC2024].
The ST programming guide presents a program creation procedure that explicitly ties label definition, ST authoring, and conversion (compilation) into an executable sequence program, including correction steps when compile errors occur [MEC2022]. Mitsubishi Electric instructions are exposed as functions and function blocks callable from ST [MEC2023, MEC2024b].
Accordingly, the project scope is a specific copilot that integrates with GX Works3 workflows: it uses directed retrieval over official documentation and internal code patterns to emit code that compiles in GX Works3. Moreover, the copilot validates outputs via the platform’s compilation and syntax-check mechanisms [MEC2022, MEC2024].
Finally, under low-data constraints, synthetic and customer-derived task variants are iteratively refined using compiler feedback, aligning with evidence that compilation-aware feedback loops yield measurable improvements in ST code generation quality [Haag2024].
@@
Figure 2: System architecture for Mitsubishi Electric ST coding assistant with RAG, competitive multi-model orchestration, and compiler-driven iterative repair.
@@
The knowledge base consolidates three retrieval segments indexed with metadata filtering: (1) function block definitions for iQ-R CPUs, augmented with instruction suffix semantics (_P for edge execution, _U for unsigned variants, EN/ENO conventions); (2) specification excerpts from Mitsubishi Electric manuals encoding syntax rules, reserved words, and Program Organization Unit (POU) structure constraints [MEC2023, MEC2024]; and (3) auxiliary context including user-contributed examples and project-specific patterns. Documents are embedded via semantic text representations using OpenAI’s text-embedding-3-large model [OpenAI2024] and retrieved through specialized query strategies tuned for definitions, rule constraints, and usage examples respectively.
@@
In contrast to Koziolek et al. [Koziolek2024b], who demonstrated retrieval-augmented generation for general IEC 61131-3 with manual IDE import and OpenPLC validation, this architecture addresses Mitsubishi Electric-specific constraints through fine-grained retrieval segmentation (libraries/specs/general vs. unified vector store), competitive multi-model orchestration (three concurrent paths vs. single GPT-4), and automated compiler-driven repair loops (bounded iteration vs. manual import and spot testing). While Koziolek et al. focused on proof-of-concept with OSCAT open-source libraries, this work targets proprietary vendor ecosystems under low-data conditions via strategic prompt engineering rather than exhaustive document chunking.
Unlike approaches requiring exclusive cloud-dependent processing [Yang2024], extensive fine-tuning and generating datasets [Haag2024, Ren2025], or single-model architectures [Liu2024], this system operates effectively in low-data proprietary environments through directed retrieval, strict vendor-specific guardrails, and optional edge deployment.
@@ III Implementation Challenges and Solutions
GX Works3’s label-based variable management diverges standard code gerneration practices, requiring external label registration rather than inline VAR declarations [MEC2024]. Generated code must reference pre-defined device labels while avoiding self-contained variable blocks that the GX Works3 converter rejects. The prompt explicitly instructs models to assume label pre-registration and provides templates demonstrating compliant device referencing patterns. We set up an MCP server that parses VAR...END_VAR logic and imports it into GX Works 3.
@@
Industrial environments often prohibit cloud dependencies due to data protection regulations [Brehme2025], necessitating on-premises inference. The local model employs quantized GGUF [ggml2025] format (deepseek-coder-v2-lite fine-tuned) [DeepSeekAI2024] to enable deployment on resource-constrained hardware including standard engineering laptops. Quantization reduces model size and memory footprint while preserving sufficient code generation capability for Mitsubishi Electric-specific tasks, balancing privacy requirements against computational limitations (here 10GB VRAM are sufficient).
@@
These challenges underscore the viability of prompt engineering combined with strategic validation as an adaptation pathway for low-data proprietary domains. By encoding vendor-specific constraints directly into prompts, augmenting documentation with machine-generated semantic annotations, and validating synthetic data through actual compilation, the system achieves practical code generation without large-scale dataset collection or extensive model retraining [Haag2024, Ren2025]. The competitive evaluation framework further enables evidence-based deployment decisions tailored to specific industrial security and performance requirements.
@@ IV Evaluation in Mitsubishi Electric Environment
With a focus on real-world applications, we incorporated extensive human feedback by releasing an early demo app for our industry partners. However, to provide a structured and empirical evaluation, we generated 100 varied queries that were used to generate 100 ST programs with each model composition. We have evaluated the four earlier mentioned models – Azure GPT-4.1 RAG — Standard (i.e., non-RAG); Local Fine-tuned RAG — Standard (i.e., DeepSeek Coder v2 Lite Instruct) – and further non-fine-tuned models for a broader view on our RAG approach. We got access to GPT-5 after finishing the project, but we also included it in our evaluation experiments (using reasoning set to low). All RAG configurations employ OpenAI’s text-embedding-3-large [OpenAI2024] for retrieval, ensuring consistent semantic search across model variants. The results can be found in Tab. I.
We also generated queries to get training data: Our local model was trained with generated ST programs that were compilable by GX Works3. The generated code was produced by the earlier described RAG system and its sophisticated prompt. To achieve a high variability, we set up different personas for the AI to match real-world query styles. Personas were, e.g., ”casual developer”, ”electrician jargon”, ”control engineer”, etc.. We added additional functional flags to make sure we have a broad variety in program categories: ”timers”, ”communication”, ”array processing”, ”pid control”, etc. Since the compiler got us a hard check, we aimed at a high number of generated programs of which only the best (i.e., those that passed) were saved for training.
Our quantitative evaluation results can be found in Tab. I. As can be seen, the local fine-tuned RAG model performed best (DeepSeek Coder v2 Lite Instruct). The second best is the fine-tuned local model without RAG that achieved a 73 % compile rate compared to 86% for the RAG. Interesting is the fast inference time and high quality of the local standard model, i.e. the one that did not use RAG. We also tested earlier versions of the DeepSeek model and others like CodeLlama [Roziere2023] and Qwen-3 Coder [Yang2025], which performed less well. Our fine-tuned DeepSeek model was able to write code in the Mitsubishi Electric ST style and use the relevant function blocks due to training. That is, it performed well with no overhead. However, an overhead was necessary to generate curated training data. The RAG version took much longer to generate code and we had to shorten the system prompt to achieve reasonable answer times (< 1 min).
@@
However, we are comparing apples with oranges here: Models like GPT-5 or even GPT-4.1, that actually followed GPT-4.5[OpenAI2025], have surely several hundred billion parameters or more, the true number is not public, while DeepSeek Coder v2 has 16 billion [GitHub.com2025]. The latter model fits into an Apple MacBooks memory, esp. when quantized, where it used about 10 GB. Hence, the users would opt for a hosted GPT-4.1/5 with RAG when data protection is not the primary concern. But, as can be seen, a local option is also available, even for highest security standards.

### v2

@@ I Introduction
Every day, millions of Programmable Logic Controllers (PLCs) orchestrate production lines, manage power grids, and safeguard critical infrastructure—from automotive assembly plants to water treatment facilities [8]. Behind each automated process lies control logic programmed in specialized languages, most prominently Structured Text (ST), a high-level language standardized in IEC 61131-3 that combines the expressiveness of modern programming with the real-time determinism industrial systems demand [8].
@@
Despite ST’s syntactic similarity to languages like Python or Pascal, developing PLC control logic remains fundamentally different. Where Python developers access thousands of open-source libraries and Stack Overflow answers, ST engineers face sparse code repositories and proprietary vendor documentation. Moreover, each manufacturer – Siemens, CODESYS, Mitsubishi Electric, Rockwell Automation – implements distinct dialects with incompatible function blocks, naming conventions, and compiler behaviors [22, 7]. A timer implementation for Siemens TIA Portal will not compile on a Mitsubishi Electric iQ-R controller; even seemingly portable constructs like counter instructions exhibit vendor-specific reset semantics and enable/disable conventions.
@@
This fragmentation collides with industrial realities: manufacturing firms guard control logic as trade secrets, regulatory frameworks restrict proprietary code sharing, and security-critical infrastructure operators prohibit cloud-dependent AI tools that could expose operational intelligence [3]. The result is a low-data domain where traditional large language model training approaches—reliant on vast public code corpora—struggle to generate vendor-compliant, compilable control logic [7]. Existing coding assistants trained on GitHub’s predominantly web and enterprise software fail to recognize that RTRIG_P triggers on rising edges in Mitsubishi Electric ST, that variable declarations must occur in external label editors rather than inline VAR blocks, or that certain ladder logic instructions are forbidden in ST contexts.
@@
Nonetheless, recent research has demonstrated significant advances in automated ST code generation using Large Language Models (LLMs) [21]. Haag et al. [7] pioneered the integration of compiler feedback with LLM training, achieving 70% compilation success rates through iterative Direct Preference Optimization (DPO). Liu et al. [11] developed Agents4PLC, achieving up to 68.8% formal verification success across 23 programming tasks. Yang et al. [22] advanced the field with AutoPLC, a framework for vendor-aware ST code generation achieving 60% compilation success across Siemens TIA Portal and CODESYS platforms. Meanwhile, Ren et al. [19] introduced MetaIndux-PLC, employing control logic-guided iterative fine-tuning.
@@
However, despite these promising academic results, significant gaps remain between research achievements and industrial deployment requirements [3]. Brehme et al.’s [3] comprehensive industry study revealed that most RAG applications in industrial settings remain in prototype stages, with data protection, security, and quality identified as paramount concerns (8.9, 8.5, and 8.7 of 10 points).
@@
Existing research has primarily focused on either general IEC 61131-3 compatibility or multiple vendor platforms simultaneously, but has not adequately addressed the challenge of creating practical coding assistants for specific proprietary implementations [3]. While Haag et al. [7] demonstrated compiler feedback integration, their approach requires extensive fine-tuning and relies on cloud-based services. Yang et al. [22] addressed multi-vendor variations but depends on cloud-based processing incompatible with industrial security requirements [3]. Liu et al. [11] operates as a single-model solution lacking comprehensive retrieval mechanisms for proprietary function blocks. Ren et al. [19] requires substantial fine-tuning datasets typically unavailable in proprietary environments.
@@
Recent work has extended LLM-based control logic generation beyond textual notations. Koziolek et al. [9] demonstrated Spec2Control, a workflow generating graphical DCS control logic for ABB systems with a high connection accuracy across 65 test cases, achieving 94-96% labor savings. While Spec2Control targets multi-plant DCS graphical programming, our work focuses on vendor-specific textual ST generation.
@@
The specialized nature of industrial automation creates a low-data domain scenario where comprehensive training datasets for specific vendor implementations are scarce or confidential [7], making traditional fine-tuning approaches impractical [3]. Notably, no existing work has specifically addressed Mitsubishi Electric’s ST implementation and MELSOFT GX Works3, the engineering environment for Mitsubishi Electric’s PLC series, which presents challenges due to its proprietary function blocks, specific syntax variations, and integration requirements. These are, however, compliant with the IEC 61131-3 standard, since it leaves freedoms to vendors. These freedoms and said function blocks, are not known to the wider public and hence not known to LLMs.
@@
This work addresses this critical gap through a novel RAG-enhanced coding assistant specifically designed for Mitsubishi Electric’s ST implementation with the development environment and compiler MELSOFT GX Works3. Unlike previous approaches targeting multiple vendors [22] or require extensive fine-tuning [7, 19], our solution focuses on one ecosystem.
@@
The key contributions include: (1) the first RAG-based coding assistant specifically designed for Mitsubishi Electric’s ST implementation, addressing a significant gap in vendor-specific industrial AI tools; (2) novel competitive multi-model approach with MELSOFT GX Works3 integration enabling real-time compilation feedback and error correction; (3) comprehensive evaluation using Mitsubishi Electric’s development environment with both compilation statistics and expert assessments from engineers familiar with the platform; and (4) practical demonstration that RAG-supported coding assistants can achieve high-quality code generation for specific proprietary platforms through targeted knowledge base construction and vendor-specific prompt engineering. We also enhance earlier work [18] that used translation transformers due to privacy constraints and plan to further support local Small and Medium Sized companies (SMEs) in the future with our results. Fig. 1 illustrates an ST program generated by our approach.
@@
Figure 1: Example of generated ST code using our GeneriST web app (built with Streamlit [20]).
@@ II Industrial Project Background and System Architecture
GX Works3 emphasizes a label- and device-based model in which variables (labels) are defined and managed in the engineering tool and then referenced by ST programs; the operating manual details dedicated editors and the standard workflow for creating ST programs, running syntax checks, and converting projects [14].
The ST programming guide presents a program creation procedure that explicitly ties label definition, ST authoring, and conversion (compilation) into an executable sequence program, including correction steps when compile errors occur [12]. Mitsubishi Electric instructions are exposed as functions and function blocks callable from ST [13, 15].
Accordingly, the project scope is a specific copilot that integrates with GX Works3 workflows: it uses directed retrieval over official documentation and internal code patterns to emit code that compiles in GX Works3. Moreover, the copilot validates outputs via the platform’s compilation and syntax-check mechanisms [12, 14].
Finally, under low-data constraints, synthetic and customer-derived task variants are iteratively refined using compiler feedback, aligning with evidence that compilation-aware feedback loops yield measurable improvements in ST code generation quality [7].
@@
Figure 2: System architecture for Mitsubishi Electric ST coding assistant with RAG, multi-model orchestration, and compiler-driven iterative repair.
@@
The knowledge base consolidates three retrieval segments indexed with metadata filtering: (1) function block definitions for iQ-R CPUs, augmented with instruction suffix semantics (_P for edge execution, _U for unsigned variants, EN/ENO conventions); (2) specification excerpts from Mitsubishi Electric manuals encoding syntax rules, reserved words, and Program Organization Unit (POU) structure constraints [13, 14]; and (3) auxiliary context including user-contributed examples and project-specific patterns. Documents are embedded via semantic text representations using OpenAI’s text-embedding-3-large model [16] and retrieved through specialized query strategies tuned for definitions, rule constraints, and usage examples respectively.
@@
In contrast to Koziolek et al. [10], who demonstrated retrieval-augmented generation for general IEC 61131-3 with manual IDE import and OpenPLC validation, this architecture addresses Mitsubishi Electric-specific constraints through fine-grained retrieval segmentation (libraries/specs/general vs. unified vector store), competitive multi-model orchestration (three concurrent paths vs. single GPT-4), and automated compiler-driven repair loops (bounded iteration vs. manual import and spot testing). While Koziolek et al. focused on proof-of-concept with OSCAT open-source libraries, this work targets proprietary vendor ecosystems under low-data conditions via strategic prompt engineering rather than exhaustive document chunking.
Unlike approaches requiring exclusive cloud-dependent processing [22], extensive fine-tuning and generating datasets [7, 19], or single-model architectures [11], this system operates effectively in low-data proprietary environments through directed retrieval, strict vendor-specific guardrails, and optional edge deployment.
@@ III Implementation Challenges and Solutions
GX Works3’s label-based variable management diverges standard code gerneration practices, requiring external label registration rather than inline VAR declarations [14]. Generated code must reference pre-defined device labels while avoiding self-contained variable blocks that the GX Works3 converter rejects. The prompt explicitly instructs models to assume label pre-registration and provides templates demonstrating compliant device referencing patterns. We set up an MCP server that parses VAR...END_VAR logic and imports it into GX Works 3.
@@
Industrial environments often prohibit cloud dependencies due to data protection regulations [3], necessitating on-premises inference. The local model employs quantized GGUF [5] format (deepseek-coder-v2-lite fine-tuned) [4] to enable deployment on resource-constrained hardware including standard engineering laptops. Quantization reduces model size and memory footprint while preserving sufficient code generation capability for Mitsubishi Electric-specific tasks, balancing privacy requirements against computational limitations (here 10GB VRAM are sufficient).
@@
These challenges underscore the viability of prompt engineering combined with strategic validation as an adaptation pathway for low-data proprietary domains. By encoding vendor-specific constraints directly into prompts, augmenting documentation with machine-generated semantic annotations, and validating synthetic data through actual compilation, the system achieves practical code generation without large-scale dataset collection or extensive model retraining [7, 19]. The competitive evaluation framework further enables evidence-based deployment decisions tailored to specific industrial security and performance requirements.
@@ IV Evaluation in Mitsubishi Electric Environment
With a focus on real-world applications, we incorporated extensive human feedback by releasing an early demo app for our industry partners. However, to provide a structured and empirical evaluation, we generated 100 varied queries that were used to generate 100 ST programs with each model composition. We have evaluated the four earlier mentioned models – Azure GPT-4.1 RAG — Standard (i.e., non-RAG); Local Fine-tuned RAG — Standard (i.e., DeepSeek Coder v2 Lite Instruct) – and further non-fine-tuned models for a broader view on our RAG approach. We got access to GPT-5 after finishing the project, but we also included it in our evaluation experiments (using reasoning set to low). All RAG configurations employ OpenAI’s text-embedding-3-large [16] for retrieval, ensuring consistent semantic search across model variants. The results can be found in Tab. I.
We generated synthetic training data for the local fine-tuned model using our RAG-enhanced pipeline with GPT-4.1. To ensure diversity reflecting real-world usage patterns, we employed multiple user personas: (1) a casual developer using informal problem descriptions, (2) an electrician employing domain-specific jargon, (3) a control engineer providing precise technical specifications. Task categories were systematically varied across timers, communication, array processing, pid control, etc. to cover common industrial automation scenarios.
Each generated sample underwent automated validation: first, GX Works3 compilation verified syntactic correctness; second, static consistency checks ensured that declared function blocks were actually invoked (rejecting samples with unused declarations) and that no forbidden identifiers or single-letter variable names appeared. Only samples passing both compilation and static validation were retained for fine-tuning. This compiler-validated filtering approach ensured training data quality without requiring manual annotation, addressing the low-data constraints inherent to proprietary industrial environments.
Our quantitative evaluation results can be found in Tab. I. As can be seen, the local fine-tuned RAG model performed best (DeepSeek Coder v2 Lite Instruct). The second best is the fine-tuned local model without RAG that achieved a 73 % compile rate compared to 86% for the RAG. Interesting is the fast inference time and high quality of the local standard model, i.e. the one that did not use RAG. We also tested earlier versions of the DeepSeek model and others like CodeLlama [2] and Qwen-3 Coder [1], which performed less well. Our fine-tuned DeepSeek model was able to write code in the Mitsubishi Electric ST style and use the relevant function blocks due to training. That is, it performed well with no overhead. However, an overhead was necessary to generate curated training data. The RAG version took much longer to generate code and we had to shorten the system prompt to achieve reasonable answer times (< 1 min).
@@
However, we are comparing apples with oranges here: Models like GPT-5 or even GPT-4.1, that actually followed GPT-4.5[17], have surely several hundred billion parameters or more, the true number is not public, while DeepSeek Coder v2 has 16 billion [6]. The latter model fits into an Apple MacBooks memory, esp. when quantized, where it used about 10 GB. Hence, the users would opt for a hosted GPT-4.1/5 with RAG when data protection is not the primary concern. But, as can be seen, a local option is also available, even for highest security standards.